# Towards Foundation Models for Experimental Readout Systems Combining Discrete and Continuous Data

### v1

@@ Abstract
We present a (proto) Foundation Model for Nuclear Physics, capable of operating on low-level detector inputs from Imaging Cherenkov Detectors at the future Electron Ion Collider. To address limitations in existing next-token prediction approaches—namely resolution loss from VQ-VAE tokenization and lack of conditional generation—we propose three key innovations: (i) separate vocabularies for discrete spatial features and continuous variates, combined via Causal Multi-Head Cross-Attention (CMHCA), (ii) continuous kinematic conditioning through prepended context embeddings, and (iii) scalable and simple, high-resolution continuous variate tokenization without joint vocabulary inflation. Our model enables fast, high-fidelity generation of pixel and time sequences for Cherenkov photons, validated through closure tests in the High Performance DIRC. We also show our model generalizes to reconstruction tasks such as pion and kaon identification, in which we show its ability to leverage fine-tuning.
May 13, 2025
Keywords: Foundation Models, Causal Cross Attention, Transformers
@@ 1 Introduction
Foundation Models (FM) are beginning to emerge as powerful tools within the field of both Nuclear Physics (NP) and High Energy Physics (HEP), supporting the inclusion of multiple downstream tasks [1, 2, 3, 4] and fine-tuning [5, 6].111In [2], the authors coin the term proto-Foundation Models, representing model development along the path to true, large-scale Foundation Models. Seminal works, such as those in [2, 3], show that relatively modest sized transformer backbones are easily generalizable to both reconstruction (particle identification (PID), jet tagging, etc) and simulation tasks. In which fine tuning from one task to another possess inherent reduction in computational overhead in comparison to models trained from scratch. Moreover, they show that different underlying mechanisms, i.e., next token prediction, masked prediction (akin to traditional decoder only, or encoder-decoder style language models) [1, 7, 6, 2, 8] or iterative refinement through Diffusion Transformers [9] can be equally as generalizable. In which the latter method [3] promotes reconstruction tasks through selective time restrictions in the diffusion processes (i.e., ). In general, both of these approaches work with relatively high-level reconstructed quantities such as 4-vectors. Recently, Birk et al. [10] demonstrated that next-token prediction can be effectively applied to simulate calorimeter response, particularly for modeling photon-induced shower development using low-level features such as pixel positions and energy depositions. While the results are promising and represent an important step forward, there are two potential limitations in the current approach that warrant further consideration. First, given the requirements of next token prediction to (generally) operate in a discrete space, quantization methods must be employed upstream in terms of a learnable tokenization process, generally in the form of a Vector Quantised-Variational AutoEncoder (VQ-VAE) [11]. VQ-VAE models provide discrete latent encodings, therefore allowing tokenization of continuous and multidimensional spaces such as location (pixel) and energy in calorimeters. While powerful, such methodologies do not come without significant tradeoffs in terms of resolutions - a potential degrading effect in physics simulations of detectors through loss of information.222Decoded tokens inherently posses some uncontrolled smearing effect proportional to the codebook size. Second, their approach does not encompass conditional generation, and therefore generative tasks that exhibit explicit conditional dependence on external parameters (e.g., Cherenkov production in DIRC detectors) is not possible. As a result, we propose two improvements for next token prediction models to circumvent these potential issues in physics, in which we deploy our techniques to Imaging Cherenkov Detectors at the future Electron Ion Collider (EIC) - pixelated detectors that require both precise location, and timing information of individual hits.333We leave translation to calorimeters (see, e.g., [10]) for future works. The main contributions of our work is as follows:
@@
The use of split vocabularies allows our model to retain high resolution time generations, without the need for excessive vocabulary size or learnable tokenization. In fact, the resolutions capable of being deployed in our study would comprise a joint-vocabulary (space and time) of roughly 36 million tokens (5920 possible time values, 6144 possible pixel values). We also note that our approximation through split vocabularies remains valid under most pixelated detector systems, given the one-to-many mappings that occur between space and the continuous variate such as time, energy or both [14]. In this scenario, one could envision the combination of space, time, and energy through sequential CMHCA blocks, where time queries energy to generate a temporal-energy informed embedding, which is then used to query spatial indices.
In what follows, we demonstrate the ability of our method to produce high fidelity fast simulation of pions and kaons in the High Performance DIRC (hpDIRC) through a series of closure tests, along with ability to operate under a classification scheme both with and without fine tuning.
@@ 2 Dataset
As stated in [15], the generated charged tracks (both pions and kaons) are distributed approximately uniformly over the phase-space, spanning GeV/c, , providing 5 million tracks for each PID. We limit ourselves to a max sequence length of 250, corresponding to a large range in photon yields across the phase-space. In contrast to prior works learning information at the photon-level, in which tracks can be formed through aggregation, we instead directly learn to generate individual tracks (sequences of hits). As a result, the transformations in [15] are not needed given that we directly work with the discretized readout (pixel indices) and time.
@@ 3 Methods
We first obtain time and spatial embeddings through learnable, independent projections from their vocabularies, in which we add positional embeddings along each dimension. We also embed both the momentum and angular values through independent linear projections (continuous spaces), and prepend our kinematics as initial context along both time and spatial projections, an example of which is given in Eq. 3. Each embedding dimension is set to 256 in our experiments.
@@
We then combine information through a transformer block using CMHCA, in which we obtain our Query projection (Q) from the time embeddings, and the Key (K) and Value (V) projections from the spatial embeddings.444For a given time, query the possible pixel locations of valid Cherenkov photons. The result is a spatio-temporally aware embedding to be further processed by transformer blocks utilizing traditional Multi-Head Self-Attention (MHSA) blocks [16]. Each transformer block consists of its respective attention type with 8 heads, a linear projection layer, and a standard feed forward network with 2 linear layers and GeLU activation [17]. Given the large combinatorial space of plausible tracks at a given kinematics, careful consideration must be taken within each attention head. Specifically, we must enforce the model to learn more focused attention maps given an initial context. We do not want the attention maps to simply collapse to the most probable configurations as aggregations over the entire phase-space. Therefore, we employ multiple normalization schemes, in which both pre and post norm strategies are used. We further augment this through normalization of the Q,K matrices [18], in which normalization is applied with a learnable scale factor. This makes attention scores invariant to their scale - promoting stability, increased variability in our output space and direction focused. This being a key feature given the large combinatorial space of plausible track configurations mentioned prior.
@@
In the case of generation, the final output from the transformer blocks feeds two independent next token prediction heads operating over the space and time vocabularies. Each of which are represented by single linear layers. In the case of classification, we represent our output latent distributions through a single linear layer. This layer is fed through a traditional CLS token format, in which we place this token at the start of each sequence. A diagrammatic representation of our model is shown in Fig. 2.
In our current study, we deploy two separate generative models for the two classes, pions and kaons (/K), as motivated by previous works and also their high degree of similarity as momentum increases. The latter requiring clear separation of their modes to provide high fidelity, and reliable simulation. We leave their combination under one model for future studies.
@@ 4 Analysis and Results
We follow the procedure in Giroux et al. [15], validating our fast simulation approach at the central region in momentum () through a series of histograms and ratios, in which we will also compare the learned sequence length, i.e., ability to learn correct photon yield as a function of the phase-space. Plots for additional kinematics (e.g., 3 GeV/c and 9 GeV/c) can be found in A and B. Note that while our model directly provides pixel and timing location, we convert the spatial indices into coordinates to provide more informative visualizations. This will then be followed by a closure test using classifier metrics, namely the same Kernel Density Estimation (KDE) derived from FastDIRC [19] at fixed momentum. In which we ultimately translate the fidelity of our generations into a more interpretable metric of separation power.
@@
We begin through visual comparisons of generations produced by our architecture to the ground truth of Geant4 at fixed momentum. We show generations from our model near the two ends of the bar (large, or small polar angle), along with the central region of the bar. As mentioned in [15], the former provides insight into our architectures ability to capture symmetry in ring structures at the extreme polar angles, while accounting for the inherent time shift that occurs due to path length differences. The latter provides insight into the ability of the architecture to model both direct, and indirect photons emitted from central regions of the bar, indicated through multi-modal structures in time. In general, smaller values of the polar angle provide insight into the ability to model both indirect and direct photons. However, in the central region the number of indirect and direct photons tends to be more similar and motivates our choice. While not quantitative in nature, fixed point generation comparisons are extreme stress tests of our model given its training on a fully continuous phase space. Moreover, these simple projections provide immediate cues into potential mode collapse, and deviation from correct rings structures.
@@
In the generations that follow, we employ a Nucleus Sampling [20] technique with , and a fixed temperature of . These generations schemes have shown to provide the most consistent cumulative distributions across the phase-space. It should also be noted that given the high degree of stochasticity within individual sequences, small changes in these parameters can result in significantly different outputs. In fact, the idea of “next token” prediction of Cherenkov photons is highly ambiguous, as shown by previous works in which high fidelity track-level generations can be produced through aggregations of individual photon generators [21, 15]. In these works, the authors show treating individual photons as independent of one another provides no loss of coherence. Fig. 3 depicts generations of kaons (left column of images) and pions (right column of images) for various polar angles ( top row, middle row and bottom row).
@@
From inspection of the figure, we can qualitatively assess that our model correctly captures kinematic dependencies in both space and time. Furthermore, we are able to obtain hints of geometrical effects, such as the kaleidoscopic effect [22] at angles closer to the expansion volumes (e.g., ). Prior fast simulation methods have been unable to capture such details. This will be discussed in more detail in later sections. To provide a more qualitative assessment of our generations, we provide cumulative distributions in Fig. 4, for (a) kaons and (b) pions.
@@
We note exceptional agreement in the spatial dimensions, in which our ratios are 1 across all values. Our time distribution is in high agreement with the ground truth along the main population (most Cherenkov hits are within the first ), although the tails of the distribution tend to struggle due to low density. This is an expected outcome in the context of modern generative models in general.
@@
Unlike previous fast simulation methods for Cherenkov detectors, in which post-hoc modeling of the photon yield must be employed, our method directly learns this variability in sequence length as a function of the phase-space. Fig. 5 depicts a comparison between the generated and ground truth photon yields at 6 GeV/c, in bins of the polar angle for (a) kaons and (b) pions.
@@
We again follow the methods devised in [15], translating fidelity measurements to a more interpretable metric of separation power using the KDE method of FastDIRC.555More information on the exact formulation of the metric, and underlying process can be found in [15]. This evaluation is performed at fixed momenta—specifically 3 GeV/c and 6 GeV/c—to emphasize fidelity testing at specific kinematic points, which serves as a stringent stress test for our generative models trained over a fully continuous kinematic space. Fig. 6 depicts the performance comparison for (top) and (bottom), for various values of the polar angle (in bins of ), with fixed sized reference populations of 800k, for Geant4, NF, and our proposed method (indicated in legend).
From inspection of the figure, we initially see that our proposed method does degrade in performance in comparison to both Geant4, and previously proposed methods such as NF. As pointed out in [15], any smoothing of the underlying probability distribution produced by our model will incur significant performance decreases in comparison to Geant4 given the ring structures only differ by a few pixels spatially at higher momentum (eventually converging at large enough momentum). In contrast to NF, the smoothing effect incurred is not inherent to the method, but rather an artifact of the data itself. Specifically, there is approximately uniform uncertainty at each forward step in time - translating to a highly ambiguous space of “next token” prediction, i.e., relatively flat attention scores. Given the inherent uncertainty in our data as to which is the most optimal next hit, different sampling techniques possess the ability to greatly alter our generations - arguably more than other AR models operating in the traditional domain of language. In light of this, we observe that post-hoc optimization can be employed to further curate generated samples. While not specifically fine tuning weights, optimizing generation parameters (e.g., temperature, dynamic temperature, values, etc) can improve fidelity and is left for future studies.666We have implemented and tested a wide variety of sampling methods, see https://github.com/wmdataphys/FM4DIRC for more details.
@@
As discussed earlier, given that our model operates in a discrete space, we are able to capture geometric effects (degree varies with angle and sampling method) such as the kaleidoscopic effect. The kaleidoscopic effect manifests as appeared “pixel preferences” in the readout system, with the effect being more pronounced the closer a track hits to the expansion volume. Fig. 7 depicts generations at the closest possible polar angle to the expansion volume ( = ), for both (a) NF, and (b) our AR approach in the figure referenced as ‘Transformer’.
@@
Particle Identification and Fine-tuning
@@
As shown in previous works [2, 3], the flexible structure of transformer backbones supports various downstream tasks (primarily fast simulation and reconstruction tasks such as PID). In what follows, we show the performance of our model, in which we fine-tune our backbone structure (essentially all transformer blocks) from a generative procedure, to one of sequence-level classification. Specifically, we classify if charged tracks originate from pions or kaons within our detector. We also show that fine-tuning our models from generation to classification provides inherent speed ups in convergence, along with slight performance increases. Fig. 8 (a) shows the performance of our model in terms of separation power at 3 Gev/c (top), and 6 Gev/c (bottom) and various values of the polar angle. We also provide performance comparisons using the NF models from [15], and the DLL method devised in [21]. The average separation power over the polar angle is reported in the legends. Fig. 8 (b) shows the accuracy of our model integrated over the phase-space as a function of training iteration (number of batches), for both fine-tuned and models trained from scratch.
@@
From inspection of Fig. 8 (a) we conclude that both NF, and our FM are able to surpass required separation power of at , as outlined in the EIC Yellow Report [23]. We also note the unique differences in their performance, in which our FM is more heavily impacted by changes in polar angle as depicted by higher non-uniformity (e.g., see separation power at 3 GeV/c). A potential artifact of model configuration, or perhaps method of kinematic conditioning. With inspection of Fig. 8 (b) we show our model is capable of fine-tuning, it benefits in terms of both computational efficiency (rate of convergence), and slightly increases in accuracy. We also note the algorithm’s clear ability to separate charged pion and kaons at 3 Gev/c, in which our separation power essentially doubles in contrast to NF. While impressive, the usefulness of such separation power at low momentum is debatable and preferred elsewhere at higher momentum where the particle identification becomes more challenging. In future studies we aim to shift some of the focus the model puts on this momentum region through momentum weighted training schemes, or trainings limited to higher impact kinematic ranges to allow increased performance.

### v2

@@ Abstract
We present a (proto) Foundation Model for Nuclear Physics, capable of operating on low-level detector inputs from Imaging Cherenkov Detectors at the future Electron Ion Collider. Building upon established next-token prediction approaches, we aim to address potential challenges such as resolution loss from existing tokenization schemes and limited support for conditional generation. We propose four key innovations: (i) separate vocabularies for discrete and continuous variates, combined via Causal Multi-Head Cross-Attention (CMHCA), (ii) continuous kinematic conditioning through prepended context embeddings, (iii) scalable and simple, high-resolution continuous variate tokenization without joint vocabulary inflation, and (iv) class conditional generation through a Mixture of Experts. Our model enables fast, high-fidelity generation of pixel and time sequences for Cherenkov photons, validated through closure tests in the High Performance DIRC. We also show our model generalizes to reconstruction tasks such as pion/kaon identification, and noise filtering, in which we show its ability to leverage fine-tuning under specific objectives.
July 18, 2025
Keywords: Foundation Models, Causal Cross Attention, Transformers, Cherenkov
@@ 1 Introduction
Foundation Models (FM) are beginning to emerge as powerful tools within the field of both Nuclear Physics (NP) and High Energy Physics (HEP), supporting the inclusion of multiple downstream tasks [1, 2, 3, 4, 5] and fine-tuning [6, 7, 8, 9].111In [2], the authors coin the term proto-Foundation Models, representing model development along the path to true, large-scale Foundation Models. Seminal works, such as those in [2, 3], show that relatively modest sized transformer backbones are easily generalizable to both reconstruction (particle identification (PID), jet tagging, etc) and simulation tasks. In which fine tuning from one task to another possess inherent reduction in computational overhead in comparison to models trained from scratch. Moreover, they show that different underlying mechanisms, i.e., next token prediction, masked prediction (akin to traditional decoder only, or encoder-decoder style language models) [1, 10, 7, 2, 11] or iterative refinement through Diffusion Transformers [12] can be equally as generalizable. In which the latter method [3] promotes reconstruction tasks through selective time restrictions in the diffusion processes (i.e., ). In general, both of these approaches work with relatively high-level reconstructed quantities such as 4-vectors. Recently, Birk et al. [13] demonstrated that next-token prediction can be effectively applied to simulate calorimeter response, particularly for modeling photon-induced shower development using low-level features such as pixel positions and energy depositions. While the results are promising and represent an important step forward, there are two potential limitations in the current approach that warrant further consideration. First, given the requirements of next token prediction to (generally) operate in a discrete space, quantization methods must be employed upstream in terms of a learnable tokenization process, generally in the form of a Vector Quantized-Variational AutoEncoder (VQ-VAE) [14]. VQ-VAE models provide discrete latent encodings, therefore allowing tokenization of continuous and multidimensional spaces such as location (pixel) and energy in calorimeters. While powerful, such methodologies do not come without significant tradeoffs in terms of resolutions - a potential degrading effect in physics simulations of detectors through loss of information.222Decoded tokens inherently posses some uncontrolled smearing effect proportional to the codebook size. Second, their approach does not encompass conditional generation, and therefore generative tasks that exhibit explicit conditional dependence on external parameters (e.g., Cherenkov production in DIRC detectors) is not possible. As a result, we propose two improvements for next token prediction models to circumvent these potential issues in physics, in which we deploy our techniques to Imaging Cherenkov Detectors at the future Electron Ion Collider (EIC) - pixelated detectors that require both precise location, and timing information of individual hits.333We leave translation to calorimeters (see, e.g., [13]) for future works. The main contributions of this work are as follows:
@@
The use of split vocabularies allows our model to retain high resolution time generations, without the need for excessive vocabulary size or learnable tokenization. In fact, the resolutions capable of being deployed in our study would comprise a joint-vocabulary (space and time) of roughly 36 million tokens (5920 possible time values, 6144 possible pixel values). We also note that our approximation through split vocabularies remains valid under most pixelated detector systems, given the one-to-many mappings that occur between space and the continuous variate such as time, energy or both [19]. In this scenario, one could envision the combination of space, time, and energy through sequential CMHCA blocks, where, e.g., time queries energy to generate a temporal-energy informed embedding, which is then used to query spatial indices.
In what follows, we demonstrate the ability of our methods to produce high fidelity fast simulation of pions and kaons in the High Performance DIRC (hpDIRC) through a series of closure tests, in which we compare generations from class-wise independent models with those stemming from a singular model with multiple experts. We also show the model’s ability to operate under a classification scheme, both at the sequence (PID) and token (noise filtering) levels, with and without fine tuning.
@@ 2 Dataset
As stated in [18], the generated charged tracks (both pions and kaons) are distributed approximately uniformly over the phase-space, spanning GeV/c, , providing 5 million tracks for each particle type. We limit ourselves to a max sequence length of 250, corresponding to a large range in photon yields across the phase-space. In contrast to prior works learning information at the photon-level, in which tracks can be formed through aggregation, we instead directly learn to generate individual tracks (sequences of hits).
@@ 3 Methods
We first obtain time and spatial embeddings through learnable, independent projections from their vocabularies, in which we add positional embeddings along each dimension. We also embed both the momentum and angular values through independent linear projections (continuous spaces), and prepend our kinematics as initial context along both time and spatial projections, an example of which is given in Eq. 3. Each embedding dimension is set to 256 in our experiments. We therefore define the following sequences:444Here, SOS and EOS denote the Start of Sequence and End of Sequence tokens, respectively.
@@
We then combine information through a transformer block using CMHCA, in which we obtain our Query projection (Q) from the time embeddings, and the Key (K) and Value (V) projections from the spatial embeddings.555For a given time, query the possible pixel locations of valid Cherenkov photons. The result is a spatio-temporally aware embedding to be further processed by transformer blocks utilizing traditional Multi-Head Self-Attention (MHSA) blocks [20]. Each transformer block consists of its respective attention type with 8 heads, a linear projection layer, and a standard feed forward neural network (FFNN) with 2 linear layers and GeLU activation [21]. Given the large combinatorial space of plausible tracks at a given kinematics, careful consideration must be taken within each attention head. Specifically, we must enforce the model to learn more focused attention maps given an initial context. We do not want the attention maps to simply collapse to the most probable configurations as aggregations over the entire phase-space. Therefore, we employ a pre-normalization scheme in the transformer blocks [22, 23]. We further augment this through normalization of the Q,K matrices [24], in which normalization is applied with a learnable scale factor. This makes attention scores invariant to their scale - promoting stability, increased variability in our output space and direction focused. This being a key feature given the large combinatorial space of plausible track configurations mentioned prior.
@@
In the case of generation, the final output from the transformer blocks feeds two independent next token prediction heads operating over the space and time vocabularies. Each of which are represented by single linear layers. The loss is then the linear combination of the Cross Entropy (CE) over both vocabularies. In the case of classification (sequence level prediction), we represent our output latent distributions through a single linear layer. This layer is fed through a traditional CLS token format, in which we place this token at the start of each sequence. The loss is then given by traditional Binary Cross Entropy (BCE). Similarly, for filtering tasks (token level prediction), we represent the per-token latent distributions through a single linear layer operating element-wise. Given that the the signal to noise ratio is high, the class of noise is generally underrepresented. As a result, we use a focal-loss [25] with and . A diagrammatic representation of our model is shown in Fig. 2.
Class Conditional Generation with a Mixture of Experts and Fixed Routing
A Mixture of Experts [26] replaces the traditional FFNN layers of a Transformer with a set of experts—multiple FFNNs designed to specialize on different aspects of data and potentially increase interpretability [27]. Rather than all tokens being passed through the same shared FFNN, each token is dynamically routed to a subset of experts based on a gating mechanism, often referred to as a Router. Let and denote the router (a simple linear matrix operation), and the th expert. The output of a MoE layer is then given by the weighted sum of the router and the experts, Eq. 2, where is the softmax of the th expert.666In Eq. (2), and stand for the sequence length and the embedding dimension, respectively.
To increase computational efficiency, Shazeer et al. [26] employ a Top- selection criteria, sparsifying the space of experts by allowing only the most probable experts to contribute. In our work, we instead deploy a class conditional selection criteria, closely resembling that of Top-k selection (see Fig. 1 in [26] for a visual depiction of Top-k selection). Specifically, given an input sequence and its associated class label (e.g., pion or kaon), we fix the routing to a subset of experts through masking of all other expert contributions. Each class (particle type) is then routed through its own subset of experts, allowing class-specific generation while sharing the bulk of the Transformer layers, i.e., simulation of pions and kaons through a unified model. This approach mimics the sparsity of Top- while introducing explicit expert-class alignment.
We explore configurations with both two, and four total experts. In the two expert case, where each class is exclusively routed through a singular expert, no load balancing is required. However, in the four (or more) expert case, each class is routed to experts where output for each class is the weighted sum of the remaining experts. This introduces the potential for expert collapse, where only one expert dominates. To mitigate this, we include a load balancing loss that encourages the routing probabilities to be approximately uniform across the selected experts, as shown in Eq. 3:
For other tasks, such as sequence (PID) or token (filtering) level classification, MoE is not used by default. In what follows we show that such a model can still be used for fine-tuning by setting the weights of the standard FFNN as the average of the experts.
@@ 4 Analysis and Results
We follow the procedure in Giroux et al. [18], validating our fast simulation approach at the central region in momentum () through a series of histograms and ratios, in which we will also compare the learned sequence length, i.e., ability to learn correct photon yield as a function of the phase-space. Plots for additional kinematics (e.g., 3 GeV/c and 9 GeV/c) can be found in A and B. Note that while our model directly provides pixel and timing location, we convert the spatial indices into coordinates to provide more informative visualizations. This will then be followed by a closure test using classifier metrics, namely the same Kernel Density Estimation (KDE) derived from FastDIRC [28] at fixed momentum. In which we ultimately translate the fidelity of our generations into a more interpretable metric of separation power. We also conduct closure tests of generative models both with (class conditional generation), and without a MoE, in which their performance should be equivalent given effective class conditioning. In the case of the former, we compare models with four and two experts, resulting in two or a singular expert per class.
@@
We begin through visual comparisons of generations produced by our architecture to the ground truth of Geant4 at fixed momentum. We show generations from our model near the two ends of the bar (large, or small polar angle), along with the central region of the bar. As noted in [18], charged particles entering the bar at polar angles near the expansion volume [29] produce a clear temporal separation between direct and indirect photons. In this case, indirect photons travel the full bar length, reflect from the far end, and return to the readout with significantly delayed arrival times. For particles traversing the central region of the bar, indirect photons reflect after approximately half the bar length, resulting in intermediate delays. Conversely, particles near the far end of the bar generate both direct and indirect photons that arrive nearly simultaneously, making their separation more difficult. These variations in photon propagation paths across polar angles result in a range of timing topologies that serve as a rigorous test of the model’s ability to capture geometric and temporal features in Cherenkov photon detection. While not quantitative in nature, fixed point generation comparisons are extreme stress tests of our model given its training on a fully continuous phase space. By forcing the model to generate many samples at isolated points it may have rarely—or even never—seen during training, we directly test its ability to generalize smoothly and coherently across the conditioning space. Moreover, these simple projections provide immediate cues into potential mode collapse or blending of kinematic dependencies, ultimately resulting in deviation from correct rings structures.
@@
In the generations that follow, we employ a Nucleus Sampling [30] technique with , and a fixed temperature of . These generations schemes have shown to provide the most consistent cumulative distributions across the phase-space. It should also be noted that given the high degree of stochasticity within individual sequences, small changes in these parameters can result in significantly different outputs. In fact, the idea of “next token” prediction of Cherenkov photons is highly ambiguous, as shown by previous works in which high fidelity track-level generations can be produced through aggregations of individually generated photons [17, 18]. These works demonstrate that modeling individual photons as independent entities does not result in a loss of coherence in the generated distributions. Fig. 3 shows simulated kaon (left column) and pion (right column) events at three representative polar angles: (top row), (middle row), and (bottom row), generated using a mixture of four experts.
@@
From inspection of the figure, we can qualitatively assess that our model correctly captures kinematic dependencies in both space and time.777In this work, all models—including those incorporating an MoE formulation—produce generations that are visually comparable. Furthermore, we are able to obtain hints of geometrical effects, such as the kaleidoscopic effect [31] at angles closer to the expansion volumes (e.g., ). To provide a more quantitative assessment of our generations, we provide cumulative distributions in Fig. 4, for kaons (left column) and pions (right column) for independent models (top row), a mixture of two experts (middle row), and a mixture of 4 experts (bottom row).
@@
We note exceptional agreement in the spatial dimensions, in which our ratios are 1 across all values. Our time distribution closely matches the ground truth over the main photon population, with most Cherenkov hits occurring within the first . Discrepancies appear beyond , where photon densities fall well below ; as a result, their impact on the overall distribution is negligible. This is an expected outcome in the context of modern generative models in general.
@@
Unlike previous fast simulation methods for Cherenkov detectors, in which post-hoc modeling of the photon yield must be employed, our method directly learns this variability in sequence length as a function of the phase-space. Fig. 5 depicts a comparison between the generated and ground truth photon yields at 6 GeV/c, in bins of the polar angle for kaons (left column) and pions (right column) for independent models (top row), a mixture of two experts (middle row), and a mixture of 4 experts (bottom row).
@@
We again follow the methods devised in [18], translating fidelity measurements to a more interpretable metric of separation power using the KDE method of FastDIRC.888The exact metric formulation and associated methodology are described in [18]. This evaluation is performed at fixed momenta—specifically 3 GeV/c and 6 GeV/c—to emphasize fidelity testing at specific kinematic points, which serves as a stringent stress test for our generative models trained over a fully continuous kinematic space, as motivated in the previous section. Fig. 6 depicts the performance comparison for (top) and (bottom), for various values of the polar angle (in bins of ), with fixed sized reference populations of 800k, for Geant4, Normalizing Flow (NF) [18], and our proposed methods of independent models, and models with two or four experts (indicated in legend).
From inspection of the figure, we initially see that our proposed method does degrade in performance in comparison to both Geant4, and previously proposed methods such as NF. As pointed out in [18], any smoothing of the underlying probability distribution produced by our model will incur significant performance decreases in comparison to Geant4 given the ring structures only differ by a few pixels spatially at higher momentum (eventually converging at large enough momentum). Given the inherent uncertainty in our data as to which is the most optimal next hit, different sampling techniques possess the ability to greatly alter our generations—arguably more than other AR models operating in the traditional domain of language. In light of this, we observe that post-hoc optimization can be employed to further curate generated samples. While not specifically fine tuning weights, optimizing generation parameters (e.g., temperature, dynamic temperature, values, etc) can improve fidelity and is left for future studies.999We have implemented and tested a wide variety of sampling methods, see https://github.com/wmdataphys/FM4DIRC for more details. We also intend to investigate increases in performance through the scaling potential of transformers, i.e., scaling both model and training data size has shown to lead to “emergent” characteristics in the space of modern FMs, potentially translating to improved performance in our case. We also observe strong agreement between independently trained generative models and those employing class-conditional generation via the MoE framework, validating the ability to simulate both pions and kaons within a single architecture.
@@
As discussed earlier, the discrete nature of our model allows it to capture geometric effects—such as the kaleidoscopic effect—whose magnitude varies with polar angle and sampling method. The kaleidoscopic effect manifests as pixel-level structure in the readout, with the effect being more pronounced the closer a track hits to the expansion volume. Fig. 7 depicts generations at the closest possible polar angle to the expansion volume ( = ), for both (a) NF, and (b) our AR approach in the figure referenced as ‘Autoregressive’.
@@
Downstream Tasks and Fine-tuning
As shown in previous works [2, 3], the flexible structure of transformer backbones supports various downstream tasks (primarily fast simulation and reconstruction tasks such as PID). In the following, we evaluate our model’s performance after fine-tuning the backbone architecture—consisting primarily of transformer blocks—from a generative task to sequence-level classification. We also introduce another downstream task in the form of noise filtering—a potentially crucial application in high rate environments at the future EIC (e.g., these studies can be of particular relevance for a sub-detector such as the dual-RICH in ePIC, where SiPM dark current rates can be as high as kHz/channel after years of radiation damage [33]). In this case, we show some inherent limitations of fine-tuning and discuss the results in the context of prior objectives. In all cases, the models are trained under identical conditions (e.g., same learning rates, optimizers, etc.), with the only difference being in their initialization strategies.
Particle Identification
@@
We classify if charged tracks (sequence level) originate from pions or kaons within our detector, and show that fine-tuning our models from generation to classification provides inherent speed ups in convergence, along with slight performance increases. In the case of class conditional models using a MoE, we show that fine-tuning still remains valid in which we initialize the weights of the standard FFNNs as the average of the experts. While it is possible to directly use a MoE within models performing reconstruction tasks, our initial performance estimates show that the additional computation is not needed and this is left for future studies. Fig. 8 (a) shows the performance of our model in terms of separation power (the number of standard deviations separating the two latent distributions) at 3 Gev/c (top), and 6 Gev/c (bottom) at various values of the polar angle. We also provide performance comparisons using the NF models from [18], and the DLL method devised in [17]. The average separation power over the polar angle is reported in the legends. Fig. 8 (b) shows the accuracy of our model integrated over the phase-space as a function of training iteration (number of batches), for both fine-tuned and models trained from scratch. Fine-tuned refers to models initialized with the weights of an independently trained generative model (e.g., one trained solely on pions or kaons), while from MoE refers to models initialized from a multi-expert (four experts) architecture.
@@
From inspection of Fig. 8 (a) we conclude that both NF, and our FM are able to surpass required separation power of at , as outlined in the EIC Yellow Report [32]. We also note the unique differences in their performance, in which our FM is more heavily impacted by changes in polar angle as depicted by higher non-uniformity (e.g., see separation power at 3 GeV/c). A potential artifact of model configuration, or perhaps method of kinematic conditioning that will be further investigated in future studies. With inspection of Fig. 8 (b) we show our model is capable of fine-tuning, it benefits in terms of both computational efficiency (rate of convergence), and slightly increases in accuracy. We also note the algorithm’s clear ability to separate charged pion and kaons at 3 Gev/c, in which our separation power essentially doubles in contrast to NF. While impressive, the usefulness of such separation power at low momentum is debatable and preferred elsewhere at higher momentum where the particle identification becomes more challenging. In future studies we aim to shift some of the focus the model puts on this momentum region through momentum weighted training schemes, or trainings limited to higher impact kinematic ranges to allow increased performance.
Noise Filtering
We classify if individual hits (token level) created by an incoming charged track are signal or noise. The model is trained irrespective of class label, i.e., both pions and kaons are included under the same model. We use a simulated dark rate in the PMTs of in which we sample randomly during training. For a given track, both the configuration (ordering in time), and the number of dark photons is stochastic, forcing our architecture to learn contextual representations of noise and signal as opposed to relying on positional information. During evaluation, we perform the same sampling procedure in which we create fifty unique representation of each track in the testing set to obtain uncertainty estimates. Fig. 9 shows the performance of our method for (a) kaons and (b) pions in terms of precision-recall curves, and noise rejection as a function of signal efficiency. The former being agnostic to the inherent class imbalance.101010Our configuration of dark rate provides approximately of total hits as noise.
From inspection of the precision-recall curves in Fig. 9, we see our model is able to efficiently able to characterize detector noise from that of signal at the token level, noting an average precision (AP) of for both pions and kaons. We also note exceptional performance in terms of noise rejection as a function of signal retention, in which the area under the curve (AUC) is . Plots for additional kinematics (e.g., 3 GeV/c and 9 GeV/c) can be found in C in which we see uniform performance at both lower and high momenta.
As mentioned prior, there exists limitations of fine-tuning the pre-existing models to downstream tasks. The main influence in the effectiveness of fine-tuning is the degree of contextual similarity between the original task and the target task. For example, in generative modeling, transformer blocks typically learn a more global context to generate coherent outputs across an entire sequence. This global understanding is highly beneficial for both generation and sequence level classification tasks but can pose challenges when adapting the same model for token level classification. Token-level classification requires a shift in focus from global to local context, where the model needs to make decisions based on finer, more immediate surroundings of each token rather than broad sequence-wide patterns. During fine-tuning for classification, the model is required to partially “unlearn” previously encoded information and instead realign it’s attention mechanisms. This contextual realignment is non-trivial and can limit the efficiency or performance of the model on the downstream task. Fig. 10 (left) shows the validation loss as a function of training iteration for both models trained from scratch and a fine-tuned generative model.111111We fine-tune from an independent generative model only, and not from a model using a MoE given the inherent limitations seen. Fig. 10 (right) shows the validation metrics, namely the precision, recall and F1 score for both fine-tuned and models trained from scratch as a function of training iteration.
From inspection of Fig. 10 it is apparent that fine-tuning poses no benefits under the task of filtering. While not detrimental in terms of performance (metrics remain similar in value), the model trained from scratch is more optimal—in which we have identified a clear limitation. In light of this, identifying task dependent behaviors is crucial as FMs become increasingly popular within the physics community. In certain scenarios, fine-tuning may fail to fully exploit the representational capacity of the pre-trained model, thereby limiting potential performance improvements.